a gun of mass 10kg fires 4 bullets per second. the mass of each bullet is 20g and the velocity of bullet when it leaves the gun 300ms .force recquired to hold the gun while firing is....
PLZ ZZZ
 

Grade:6

1 Answers

Arun
6 years ago
MASS OF THE BULLET=10 KG
AND VELOCITY OF BULLET=300 ms^-1
THEREFORE FORCE ON BULLET=10*4*300
= 12000
FORCE IS DIRECTIONALLY PROPORTIONAL TO RATE OF CHANGE OF MOMENTUM
THEREFOR PER SECOND A BULLET APPLIES A FORCE OF MV
ACCELARATION =V-U/T
U-0
T = 1
THEREFORE A-V
THEREFORE FORCE BY ONE BULLET=MV
= 20*300
= 6000

Regards
Arun (askIITians forum expert)